# Merger History of Clustered Primordial Black Holes

### v2

@@ Abstract
Primordial black hole (PBH) binaries experience strong gravitational perturbations in the case of their initial clustering, which significantly affects the dynamics of their mergers. In this work, we develop a new formalism to account for these perturbations and track the evolution of the binary orbital parameters distribution. Based on this approach, we calculate the merger rate of PBH binaries and demonstrate that its temporal evolution differs greatly from that of isolated binary systems. Moreover, PBH clustering produces distinctive features in the stochastic gravitational-wave background: the canonical 2/32/32 / 3 spectral slope transforms to $\Omega_{\rm gw}\propto\nu^{-65/28}$ in a certain frequency band. These predictions can be probed in future gravitational wave observations, opening up new opportunities to test the clustering of PBHs and their contribution to dark matter.
@@ I Introduction
Progress in gravitational-wave astronomy has led to a surge of interest in primordial black holes (PBHs) — hypothetical black holes that may have formed in the very early Universe. Zel’dovich and Novikov (1967); Hawking (1971). In recent years, the LIGO-Virgo-KAGRA collaboration has detected dozens of compact object merger events Abbott et al. (2023a, b). Most of them are interpreted as mergers of black holes of astrophysical origin, but the primordial nature of a part of the detected events is also popular Bird et al. (2016); Sasaki et al. (2016); Clesse and Garc´ıa-Bellido (2017a); Blinnikov et al. (2016); Ali-Ha¨ımoud et al. (2017); Raidal et al. (2019); Liu et al. (2019); De Luca et al. (2020); Hütsi et al. (2021). This interpretation is particularly valuable because it allows us to use gravitational-wave (GW) observations as a tool to impose constraints on the density of PBHs in the Universe. Moreover, many other independent observations have imposed additional constraints, but primordial black holes remain a viable candidate for dark matter Carr et al. (2021) — one of the central problems of modern fundamental science.
@@
It is traditionally considered that PBH binaries are formed at the radiation-dominated stage of the Universe evolution due to fluctuations in their spatial distribution Nakamura et al. (1997); Sasaki et al. (2016). Despite a well theoretical elaboration of this scenario, quantitative estimates of the merger rate face uncertainties. The main difficulties are related to various factors that affect the binary systems: in particular, the interaction of the binaries with the dark matter halo that forms around them Kavanagh et al. (2018); Pilipenko et al. (2022); Jangra et al. (2023); Stasenko and Eroshenko (2024); perturbations of black hole pairs in early structures arising from their natural Poisson clustering Raidal et al. (2019); Vaskonen and Veermäe (2020); Jedamzik (2020); Tkachev et al. (2020); Stasenko and Belotsky (2023); Stasenko (2024); Delos et al. (2024); and other possible effects.
@@
In this paper, we consider an alternative scenario assuming an initially clustered birth of PBHs. In this case, the statistics of PBH pairs is formally constructed similarly to the model of their random distribution in space Nakamura et al. (1997); Sasaki et al. (2016). However, in the case of strong clustering, the key difference is that the PBH binaries appear to be embedded in a dense environment. In such conditions they are subjected to continuous gravitational perturbations from the surrounding single PBHs. In this paper we develop a new formalism describing how these perturbations affect mergers of PBH binaries. Our calculations show that this process leads to a significantly different temporal evolution of the merger rate compared to the prediction for pairs of PBHs without initial clustering. In particular, intense perturbations can both destroy the existing binary systems and significantly modify the orbital parameters of the surviving pairs, increasing their lifetime. These effects have a significant impact on the spectral characteristics of the stochastic GW background, which integrates the contribution of black hole mergers over cosmological history to the GW energy density in the Universe. The results have important observational implications for next generation of gravitational-wave experiments such as LISA Amaro-Seoane et al. (2017), TianQin Luo et al. (2016), Einstein Telescope Maggiore et al. (2020); Abac et al. (2025) and other proposed projects. Their expected sensitivity will allow both the detection of individual black hole merger events at high redshifts and the measurement of the gravitational-wave background over a wide frequency range, providing a test of the PBH idea.
@@
The paper is organized as follows. Section II develops the simple theoretical formalism for PBH cluster formation. In Section III, we analyze the formation of PBH binaries during the radiation-dominated era of the Universe. Section IV provides both qualitative analysis of the perturbation processes affecting binary systems and analytical estimates for the parameters of binaries undergoing perturbations. Section V presents our calculations of the PBH merger rate incorporating these perturbation effects. The implications of our results for the stochastic gravitational wave background are discussed in Section VI. Finally, Section VII summarizes our main findings.
@@ II Primordial black hole clustering
A number of theoretical models predict that the formation of PBHs is accompanied by their strong spatial clustering, which has been shown, for example, in the works Rubin et al. (2001); Khlopov et al. (2005); Young and Byrnes (2015); Tada and Yokoyama (2015); Belotsky et al. (2019); Desjacques and Riotto (2018); Suyama and Yokoyama (2019); Ding et al. (2019). In this case, there must exist regions of some spatial scale with nonzero PBH density and the abundance of such clusters in the Universe is determined by the relative density of the PBHs $\Omega_{\rm PBH}$. To parameterize the inhomogeneous PBH distribution and account for their locally density enhancements, we introduce the density contrast $\delta_{\rm PBH}(\bf x)=\rho_{\rm PBH}(\bf x)/\rho_{\rm DM}$, where $\rho_{\rm DM}$ is the background dark matter density. The quantity $\delta_{\rm PBH}$ is model-dependent, but in this work we consider it as free parameter, the specific choice of which determines the moment of gravitational detachment of the cluster from the Hubble flow. Although the quantity $\delta_{\rm PBH}$ should generally vary across spatial scales, we ignore its scale dependence (power spectrum) since our goal is to show the influence of clustering on PBHs mergers. Therefore we assume clusters have a monochromatic mass function, and in our simplified model $\delta_{\rm PBH}$ is nonzero only within some small spatial region of comoving size $R\leq R_{\rm cl}$ from the cluster center:
@@
where $\rho_{\rm DM,0}\approx 33\,M_{\odot}$ kpc-3 is the modern dark matter density. The scales under consideration are much smaller than those observed by the Planck satellite, at which constraints are imposed on the isocurvature modes. Thus, PBH clusters make up only a small part of the dark matter bulk, and their physical size are much smaller than the distance between them $r_{\rm cl}\ll R_{\rm cl}/f^{1/3}$.
@@ III Primordial black hole binaries
During the radiation-dominated stage, the PBHs form gravitationally bound pairs. This occurs because at some moment of time $t_{\rm d}<t_{\rm eq}$ the local density of two PBHs will exceed the radiation density $\rho_{r}(t_{\rm d})$. In this sense, the situation is completely analogous to the formation of clusters discussed in the previous section. To describe this process, we introduce the mean physical distance between the PBHs at the moment $t_{\rm eq}$
@@
Since the radiation density falls as $\rho_{r}\propto s^{-4}$, the moment of gravitational detachment of the pair is determined by the following expression
$\frac{m}{R^{3}}=\rho_{r}(t_{d})=\rho_{r,\rm eq}\left(\frac{s_{{\rm eq}}}{s_{\rm d}}\right)^{4},$ (8)
here $R=xs_{d}/s_{eq}$ is the physical distance between the PBHs at the moment $t_{d}$, and $x$ is the distance at the moment $t_{\rm eq}$. Thus, the moment of formation of the binary system is determined as
$\frac{s_{\rm eq}}{s_{\rm d}}=\delta_{{\rm PBH}}\left(\frac{\overline{x}}{x}\right)^{3}.$ (9)
For the detachment moment in the described formalism, there are two natural bounds: $s_{{\rm d}}<s_{\rm eq}$ in the case $\delta_{{\rm PBH}}<1$ and $s_{{\rm d}}<s_{\rm eq}/\delta_{{\rm PBH}}$ when $\delta_{{\rm PBH}}>1$. The second condition arises because the gravitational detachment of the pair from the Hubble flow must occur before the formation of the cluster. Taking this into account, as well as the fact that $x<\overline{x}$, we conclude that pairs are formed only by those PBHs whose distance between them satisfies the relation $x<\overline{x}\delta_{\rm PBH}^{\gamma}$, where $\gamma$ is used to parameterize two possible clustering modes: $\gamma=1/3$ for $\delta_{\rm PBH}<1$ and $\gamma=0$ for the case $\delta_{\rm PBH}>1$.
After detachment from the expansion of the Universe, the PBHs avoid a head-on collision and move in an elliptical orbit around each other because tidal torques from neighboring PBHs induce angular momentum. Here, for simplicity, we consider only the main contribution from the nearest PBH, the distance to which we denote by $y$, and the condition $x<y<\overline{x}$ must be satisfied. We follow the formalism of work Sasaki et al. (2016) and take the probability distribution over $x$ and $y$ to be flat in three-dimensional space
@@
further, the coefficients $\alpha$ and $\beta$ are taken to be equal to unity. The eccentricity of a binary is defined as $e=\sqrt{1-b^{2}/a^{2}}$, but it is more convenient to use the dimensionless angular momentum $j=\sqrt{1-e^{2}}=(x/y)^{3}$ instead. To characterize the PBH binaries it is convenient to transform to the variables $a$ and $j$ in the distribution (10)
@@
where the variables vary in the range $0<a<\overline{x}\delta_{\rm PBH}^{4\gamma-1}$, $(a\delta_{\rm PBH}/\overline{x})^{3/4}<j<1$. The lower bound on $j$ corresponds to the fact that the distance to the third PBH is maximized $y=\overline{x}$. From the distribution of (12) we can see that PBH binaries mainly have small angular momenta at formation, i.e., their orbits are strongly elongated. Due to the emission of gravitational waves a pair of black holes of the same mass $m$ merge in time Peters (1964)
@@
where $t_{0}=13.8$ Gyr is the age of the Universe. Since $a_{\rm ch}$ depends very weakly on $t_{\rm mer}$, almost all merging binaries at not too far cosmological distances are characterized by approximately the same value of the large semi-axis. Now let us compare the obtained value with the maximum possible major semiaxis. In the case of $\delta_{\rm PBH}<1$, the limiting value of $a_{\rm max}$ corresponds to the average distance between the PBHs at $\delta_{\rm PBH}=1$, which is $a_{\rm max}\sim 10^{4}$ au. In the case $\delta_{\rm PBH}>1$, we get $a_{\rm max}\sim 10^{4}\delta_{\rm PBH}^{-4/3}$ au. Thus, all PBH binaries merging by the modern epoch formed gravitationally bound pairs long before the transition to the stage of matter dominance (or before the formation of the cluster $t_{f}$ at $\delta_{\rm PBH}>1$). Substituting the found expression for the semi-major axis (III) into $j_{\rm min}=(a\delta_{\rm PBH}/\overline{x})^{3/4}$ one can determine the characteristic value of the angular momentum of the binaries
@@ IV Perturbations of binaries
Binary systems with large semi-major axis $a>a_{h}$ are classified as soft, otherwise — hard binaries. According to the classical Heggi-Hill law, such hard binaries tend to become harder when interacting with single objects ($|E_{b}|$ increases), while soft ones become softer Heggie (1975); Hills (1975).
@@
here $v$ — the relative velocity between the binary and single PBH at infinity. In our simplified model it is assumed that, as a result of a single act of hard scattering, the angular momentum of the binary increases to $j\sim 1$ and hence its merger time increases significantly. To define hard scattering we use a simple criterion, which is that this distance of minimum approaching is equal to the major semiaxis $r_{p}=a$. The probability of such interactions is determined by the mean free time $\tau^{-1}=\rho_{{\rm PBH}}\langle\Sigma_{p}v\rangle/m$. Assuming the Maxwell distribution of PBHs velocities in the cluster with dispersion $\sigma$, and after averaging over the relative velocities, we obtain the characteristic perturbation time of orbital parameters of PBH binaries
@@
where the density of black holes in the cluster is $\rho_{\rm PBH}\approx 140\,\delta_{\rm PBH}^{4}\rho_{\rm eq}$ (see Section II). First of all, let us estimate the perturbation time of the characteristic binaries (III) merging in the modern epoch. As an example, we consider the case of the PBH with mass $m=10\,M_{\odot}$ with clustering parameter $\delta_{\rm PBH}=1$ and velocity dispersion $\sigma=10$ km s-1. The perturbation time is estimated to be $\tau\approx 0.6$ Myr, and since the semimajor axis of typical binaries (III) depends weakly on the merger time, it can be concluded that typical binaries are subject to perturbation for almost the entire cosmological history after their formation. Let us now consider the case of binaries in a circular orbit and with coalescence time $t_{0}$. At the time of formation, the semimajor axis of such a binary should be $a\approx 0.8$ au, and their perturbation time turns out to be $\tau\approx 2$ Gyr. Note that for a cluster with velocity dispersion $\sigma>75$ km s-1 these binaries turn out to be even soft. These estimates show that in the considered example all binaries that potentially produce observable GW signals are perturbed in clusters. In the case of earlier time scales $<t_{0}$ or with decreasing $\delta_{\rm PBH}$ and PBH masses, the perturbations affect only a certain fraction of all binaries, namely only sufficiently wide binaries are subject to perturbations.
[…]
Let us now consider binaries with some merger time $t_{\rm mer}$. Using Eq. (13), we express the major semiaxis through $t_{\rm mer}$ and angular momentum $j$ to obtain the perturbation time $\tau(t_{\rm mer},j)$. The binaries with $\tau>t_{\rm mer}$ merge before a significant perturbation, while in the opposite case at $\tau<t_{\rm mer}$ the perturbation by a third black hole leads to an increase of the binary merger time. The angular momentum defining the boundary between these regimes is found from the solution of the equation $\tau(t,j_{p})=t$:
@@
where we used the gravitational focusing approximation in the cross section (17). As can be seen from the comparison of the Eqs. (15) and (19), most of the binaries that would merge in the contemporary Universe are perturbed since $j_{p}\gg j_{\rm ch}$. Thus, a constraint on the angular momentum of merging binaries $1<j<j_{p}$ arises, as is clearly demonstrated by Fig. 1. Let us now define the moment of time $t_{\rm sup}$, starting from which the characteristic binaries are subject to perturbations, as a solution to the equation $\tau(a_{\rm ch})=t$:
@@
Alternatively, this moment can be defined as the intersection of the $j_{p}$ and $j_{\rm min}$ curves, which is also reflected in Fig. 1. For $t<t_{\rm sup}$, the merger time $t_{\rm mer}$ turns out to be shorter than the perturbation time $\tau$ even for binaries with the largest possible semimajor axis. However, for $t>t_{\rm sup}$, the merger suppression mechanism comes into play due to gravitational perturbations of the binaries by the surrounding PBHs.

### v3

@@ Abstract
Primordial black hole (PBH) binaries experience strong gravitational perturbations in the case of their initial clustering, which significantly affects the dynamics of their mergers. In this work, we develop a new formalism to account for these perturbations and track the evolution of the binary orbital parameters distribution. Based on this approach, we calculate the merger rate of PBH binaries and demonstrate that its temporal evolution differs greatly from that of isolated binary systems. Moreover, PBH clustering produces distinctive features in the stochastic gravitational-wave background: the canonical $2/3$ spectral slope transforms to $\Omega_{\rm gw}\propto\nu^{-65/28}$ in a certain frequency band. These predictions can be probed in future gravitational wave observations, opening up new opportunities to test the clustering of PBHs and their contribution to dark matter.
@@ I Introduction
Progress in gravitational-wave astronomy has led to a surge of interest in primordial black holes (PBHs) — hypothetical black holes that may have formed in the very early Universe. Zel’dovich and Novikov (1967); Hawking (1971). In recent years, the LIGO-Virgo-KAGRA collaboration has detected dozens of compact object merger events Abbott et al. (2023a, b). Most of them are interpreted as mergers of black holes of astrophysical origin, but the primordial nature of a part of the detected events is also popular Bird et al. (2016); Sasaki et al. (2016); Clesse and García-Bellido (2017a); Blinnikov et al. (2016); Ali-Haïmoud et al. (2017); Raidal et al. (2019); Liu et al. (2019); De Luca et al. (2020); Hütsi et al. (2021); Fakhry et al. (2021, 2022); Fakhry (2024). This interpretation is particularly valuable because it allows us to use gravitational-wave (GW) observations as a tool to impose constraints on the density of PBHs in the Universe. Moreover, many other independent observations have imposed additional constraints, but primordial black holes remain a viable candidate for dark matter Carr et al. (2021) — one of the central problems of present-day fundamental science.
@@
It is traditionally considered that PBH binaries are formed at the radiation-dominated stage of the Universe evolution due to fluctuations in their spatial distribution Nakamura et al. (1997); Sasaki et al. (2016). Despite a well theoretical elaboration of this scenario, quantitative estimates of the merger rate face uncertainties. The main difficulties are related to various factors that affect the binary systems: in particular, the interaction of the binaries with the dark matter halo that forms around them Kavanagh et al. (2018); Pilipenko et al. (2022); Jangra et al. (2023); Stasenko and Eroshenko (2024); perturbations of PBH binaries in early structures arising from their natural Poisson clustering Raidal et al. (2019); Vaskonen and Veermäe (2020); Jedamzik (2020); Tkachev et al. (2020); Stasenko and Belotsky (2023); Stasenko (2024); Delos et al. (2024); and other possible effects.
@@
In this paper, we consider an alternative scenario assuming an initially clustered birth of PBHs. In this case, the statistics of PBH binaries is formally constructed similarly to the model of their random distribution in space Nakamura et al. (1997); Sasaki et al. (2016). However, in the case of strong clustering, the key difference is that the PBH binaries appear to be embedded in a dense environment. In such conditions they are subjected to continuous gravitational perturbations from the surrounding single PBHs. In this paper we develop a new formalism describing how these perturbations affect mergers of PBH binaries. This aspect has not been considered in previous papers Bringmann et al. (2019); Young and Byrnes (2020); Atal et al. (2020); Crescimbeni et al. (2025) examining the impact of initial clustering on PBH mergers. Our calculations show that this process leads to a significantly different temporal evolution of the merger rate compared to the prediction for PBHs without initial clustering. In particular, intense perturbations can both destroy the existing binary systems and significantly modify the orbital parameters of the surviving binaries, increasing their lifetime. These effects have a significant impact on the spectral characteristics of the stochastic GW background, which integrates the contribution of black hole mergers over cosmological history to the GW energy density in the Universe. The results have important observational implications for next generation of GW experiments such as LISA Amaro-Seoane et al. (2017), TianQin Luo et al. (2016), Einstein Telescope Maggiore et al. (2020); Abac et al. (2025) and other proposed projects. Their expected sensitivity will allow both the detection of individual black hole merger events at high redshifts and the measurement of the GW background over a wide frequency range. This will provide a test of the PBH idea that was discussed in a series of papers, for example in Refs. Ng et al. (2022, 2023); Branchesi et al. (2023); Franciolini et al. (2023); Bagui et al. (2023); Abac et al. (2025).
@@
The paper is organized as follows. Section II develops the simple theoretical formalism for PBH cluster formation. In Section III, we analyze the formation of PBH binaries during the radiation-dominated era of the Universe. Section IV provides both qualitative analysis of the perturbation processes affecting binary systems and analytical estimates for the parameters of binaries undergoing perturbations. Section V presents our calculations of the PBH merger rate incorporating these perturbation effects. The implications of our results for the stochastic GW background are discussed in Section VI. Finally, Section VII summarizes our main findings.
@@ II Primordial black hole clustering
A number of theoretical models predict that the formation of PBHs is accompanied by their strong spatial clustering, which has been shown, for example, in the works Rubin et al. (2001); Khlopov et al. (2005); Young and Byrnes (2015); Tada and Yokoyama (2015); Belotsky et al. (2019); Desjacques and Riotto (2018); Suyama and Yokoyama (2019); Ding et al. (2019). In this case, there must exist regions of some spatial scale with nonzero PBH density and the abundance of such clusters in the Universe is determined by the relative density of the PBHs $\Omega_{\rm PBH}$. To parametrize the inhomogeneous PBH distribution and account for their locally density enhancements, we introduce the density contrast $\delta_{\rm PBH}(\bf x)=\rho_{\rm PBH}(\bf x)/\rho_{\rm DM}$, where $\rho_{\rm DM}$ is the background dark matter density. The quantity $\delta_{\rm PBH}$ is model-dependent, but in this work we consider it as free parameter, the specific choice of which determines the moment of gravitational detachment of the cluster from the Hubble flow. Although the quantity $\delta_{\rm PBH}$ should generally vary across spatial scales, we ignore its scale dependence (power spectrum) since our goal is to show the influence of clustering on PBHs mergers. Therefore we assume clusters have a monochromatic mass function, and in our simplified model $\delta_{\rm PBH}$ is nonzero only within some small spatial region of comoving size $R\leq R_{\rm cl}$ from the cluster center:
@@
where $\rho_{\rm DM,0}\approx 33\,M_{\odot}$ kpc-3 is the present-day dark matter density. The scales under consideration are much smaller than those observed by the Planck satellite, at which constraints are imposed on the isocurvature modes. Thus, PBH clusters make up only a small part of the dark matter bulk, and their physical size are much smaller than the distance between them $r_{\rm cl}\ll R_{\rm cl}/f^{1/3}$.
@@ III Primordial black hole binaries
During the radiation-dominated stage, the PBHs form gravitationally bound binary systems. This occurs because at some moment of time $t_{\rm dec}<t_{\rm eq}$ the local density of two PBHs will exceed the radiation density $\rho_{r}(t_{c})$. At the same time, these black holes decouple from the cosmological expansion and form the binary. In this sense, the situation is completely analogous to the formation of clusters discussed in the previous section. In this Section, we will assume that PBHs are uniformly distributed within the region with their increased density (in the cluster). Therefore, the results of Ref. Sasaki et al. (2016) are applicable to our consideration. First of all, we introduce the mean physical distance between the PBHs at the moment $t_{\rm eq}$
@@
Since the radiation density falls as $\rho_{r}\propto s^{-4}$, the moment of gravitational decoupling of the two PBHs is determined by the following expression
$\frac{m}{R^{3}}=\rho_{r}(t_{\rm dec})=\rho_{r,\rm eq}\left(\frac{s_{{\rm eq}}}{s_{\rm dec}}\right)^{4},$ (8)
here $R=xs_{\rm dec}/s_{{\rm eq}}$ is the physical distance between the PBHs at the moment $t_{\rm dec}$, and $x$ is the distance at the moment $t_{\rm eq}$. Thus, the moment of formation of the binary system is determined as
$\frac{s_{\rm eq}}{s_{\rm dec}}=\delta_{{\rm PBH}}\left(\frac{\overline{x}}{x}\right)^{3}.$ (9)
In the described formalism, the decoupling moment has two natural bounds: $s_{\rm dec}<s_{\rm eq}$ in the case $\delta_{{\rm PBH}}<1$ (similar unclustered scenario) and $s_{\rm dec}<s_{\rm eq}/\delta_{{\rm PBH}}$ when $\delta_{{\rm PBH}}>1$. The second condition arises because the gravitational detachment of the PBH pair from the Hubble flow must occur before the formation of the cluster. As was shown in Sec. II, the moment of PBHs cluster formation is given by $s_{f}\sim s_{\rm eq}/\delta_{\rm PBH}$, and the formation of a PBH binary must precede the formation of the cluster $s_{\rm dec}<s_{f}$, as the reverse is not physically sensible. Taking this into account, as well as the fact that $x<\overline{x}$, we conclude that binaries are formed only by those PBHs whose distance between them satisfies the relation $x<\overline{x}\delta_{\rm PBH}^{\gamma}$, where $\gamma$ is used to parametrize two possible clustering modes: $\gamma=1/3$ for $\delta_{\rm PBH}<1$ and $\gamma=0$ for the case $\delta_{\rm PBH}>1$.
After detachment from the expansion of the Universe, the PBHs avoid a head-on collision and move in an elliptical orbit around each other because tidal torques from neighboring PBHs induce angular momentum. Here, for simplicity, we consider only the main contribution from the nearest PBH, the distance to which we denote by $y$, and the condition $x<y<\overline{x}$ must be satisfied. We continue to follow the formalism of work Sasaki et al. (2016) and take the probability distribution over $x$ and $y$ to be flat in three-dimensional space
@@
further, the coefficients $\alpha$ and $\beta$ are taken to be equal to unity. The eccentricity of a binary is defined as $e=\sqrt{1-b^{2}/a^{2}}$, but it is more convenient to use the dimensionless angular momentum $j=\sqrt{1-e^{2}}=(x/y)^{3}$ instead. To characterize the PBH binaries, it is suitable to transform to the variables $a$ and $j$ in the distribution (10)
@@
where the variables vary in the range $0<a<\overline{x}\delta_{\rm PBH}^{4\gamma-1}$, $(a\delta_{\rm PBH}/\overline{x})^{3/4}<j<1$. The lower bound on $j$ corresponds to the fact that the distance to the third PBH is maximized $y=\overline{x}$. From the distribution of (12) we can see that PBH binaries mainly have small angular momenta at formation, i.e., their orbits are strongly elongated. Due to the emission of GW a binary black holes of the same mass $m$ merge in time Peters (1964)
@@
where $t_{0}=13.8$ Gyr is the age of the Universe. Since $a_{\rm ch}$ depends very weakly on $t_{\rm mer}$, almost all merging binaries at not too far cosmological distances are characterized by approximately the same value of the large semi-axis. Now let us compare the obtained value with the maximum possible major semiaxis. In the case of $\delta_{\rm PBH}<1$, the limiting value of $a_{\rm max}$ corresponds to the average distance between the PBHs at $\delta_{\rm PBH}=1$, which is $a_{\rm max}\sim 10^{4}$ au. In the case $\delta_{\rm PBH}>1$, we get $a_{\rm max}\sim 10^{4}\delta_{\rm PBH}^{-4/3}$ au. Thus, all PBH binaries merging by the current epoch formed gravitationally bound pairs long before the transition to the stage of matter dominance (or before the formation of the cluster $t_{f}$ at $\delta_{\rm PBH}>1$). Substituting the found expression for the semi-major axis (III) into $j_{\rm min}=(a\delta_{\rm PBH}/\overline{x})^{3/4}$ one can determine the characteristic value of the angular momentum of the binaries
@@ IV Perturbations of binaries
Binary systems with large semi-major axis $a>a_{h}$ are classified as soft, otherwise — hard binaries. According to the classical Heggie-Hills law, such hard binaries tend to become harder when interacting with single objects ($|E_{b}|$ increases), while soft ones become softer Heggie (1975); Hills (1975).
@@
Figure 1: The example of the orbital parameter space of PBH binaries with $\delta_{\rm PBH}=0.5$, $\sigma=10$ km s-1 and $m=10\,M_{\odot}$. Left: Binaries merging by time $\leq t_{0}/10$ are shaded in gray. The hatched area (which emphasizes that binaries with these parameters “drop out” of the merger process) schematically shows binaries that have been perturbed at the same time, for which the semimajor axis exceeds the value given by $\tau(a)=t_{0}/10$. For reasons given in the text, perturbations of binaries with a nearly circular orbit ($j\geq 1/2$) are assumed not to affect their merger time. The dashed horizontal line indicates the critical angular momentum of binaries $j_{p}$ (19), separating tight systems that avoid perturbation to $t_{0}/10$ and wide perturbed binaries. Right: Similar to the left parameter space, but on the $(t,j)$ plane, where the transformation from $a$ to $t$ is done using Eq. (13). At fixed time $t$, the merging binaries have angular momenta from $j_{\rm min}$ to 11. However, systems with small angular momenta (corresponding to wide binaries) are perturbed, leading to an effective constraint from below on the value of $j_{p}$.
@@
here $v$ is the relative velocity between the binary and single PBH at infinity. In our simplified model it is assumed that, as a result of a single act of hard scattering, the angular momentum of the binary increases to $j\sim 1$ and hence its merger time increases significantly. To define hard scattering we use a simple criterion, which is that this distance of minimum approaching is equal to the major semiaxis $r_{p}=a$. The probability of such interactions is determined by the mean free time $\tau^{-1}=\rho_{{\rm PBH}}\langle\Sigma_{p}v\rangle/m$. Assuming the Maxwell distribution of PBHs velocities in the cluster with dispersion $\sigma$, and after averaging over the relative velocities, we obtain the characteristic perturbation time of orbital parameters of PBH binaries
@@
where the density of black holes in the cluster is $\rho_{\rm PBH}\approx 140\,\delta_{\rm PBH}^{4}\rho_{\rm eq}$ (see Section II). First of all, let us estimate the perturbation time of the characteristic binaries (III) merging in the current epoch. As an example, we consider the case of the PBH with mass $m=10\,M_{\odot}$ with clustering parameter $\delta_{\rm PBH}=1$ and velocity dispersion $\sigma=10$ km s-1. The perturbation time is estimated to be $\tau\approx 0.6$ Myr, and since the semimajor axis of typical binaries (III) depends weakly on the merger time, it can be concluded that typical binaries are subject to perturbation for almost the entire cosmological history after their formation. Let us now consider the case of binaries in a circular orbit and with coalescence time $t_{0}$. At the time of formation, the semimajor axis of such a binary should be $a\approx 0.8$ au, and their perturbation time turns out to be $\tau\approx 2$ Gyr. Note that for a cluster with velocity dispersion $\sigma>75$ km s-1 these binaries turn out to be even soft. These estimates show that in the considered example all binaries that potentially produce observable GW signals are perturbed in clusters. In the case of earlier time scales $t_{0}$ or with decreasing $\delta_{\rm PBH}$ and PBH masses, the perturbations affect only a certain fraction of all binaries, namely only sufficiently wide binaries are subject to perturbations.
Let us now consider binaries with some merger time $t_{\rm mer}$. Using Eq. (13), we express the major semiaxis through $t_{\rm mer}$ and angular momentum $j$ to obtain the perturbation time $\tau(t_{\rm mer},j)$. The binaries with $\tau>t_{\rm mer}$ merge before a significant perturbation, while in the opposite case at $\tau<t_{\rm mer}$ the perturbation by a third black hole leads to an increase of the binary merger time. The angular momentum $j_{p}$ defining the boundary between these regimes is found from the solution of the equation $\tau(t,j_{p})=t$. To transform the dependence on the semi-major axis to time and angular momentum $\tau(a)\rightarrow\tau(t,j)$, the binary merger time (13) are used. After the appropriate mathematical transformations we obtain
@@
where we used the gravitational focusing approximation in the cross section (17): the second term in the brackets of Eq. (18) is the leading one. As can be seen from the comparison of the Eqs. (15) and (19), most of the binaries that would merge in the contemporary Universe are perturbed since $j_{p}\gg j_{\rm ch}$. Thus, a constraint on the angular momentum of merging binaries $1<j<j_{p}$ arises, as is clearly demonstrated by Fig. 1. Let us now define the moment of time $t_{\rm sup}$, starting from which the characteristic binaries are subject to perturbations, as a solution to the equation $\tau(a_{\rm ch})=t$:
@@
This expression is also obtained in the gravitational focusing approximation; the geometric cross-section limit for completeness is provided in Table 1. Alternatively, this moment can be defined as the intersection of the $j_{p}$ and $j_{\rm min}$ curves, which is also reflected in Fig. 1. For $t<t_{\rm sup}$, the merger time $t_{\rm mer}$ turns out to be shorter than the perturbation time $\tau$ even for binaries with the largest possible semimajor axis. However, for $t>t_{\rm sup}$, the merger suppression mechanism comes into play due to gravitational perturbations of the binaries by the surrounding PBHs.